• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 07:11:11 on April19,2004

    Entry number 118559

    This entry is a followup to: 118490

    keyword=septum PLC beam current test

    at about 0640, we asked for 75 microA to test the PLC radiation theory. we got about 70 microA. we then requested to go to 100 microA. the beam current never got above 90 microA before tripping. runs 2551 and 2552 were taken during this test. fig 1 is the PLC health. it looked like that before, during, and after the test. fig 2 is the left septum strip during that time.


    FIGURE 1

    FIGURE 2